- a charge +Q at a distance 2d from a charge -Q and a point X at a distance d from -Q what is the field strength at X numerically using SI units
1
Expert's answer
2021-10-13T09:08:35-0400

electric field at a point X:


"E=-\frac{kQ}{d^2}+\frac{kQ}{(2d+d)^2}=\frac{kQ-9kQ}{9d^2}=-\frac{8kQ}{9d^2}"